B4K HE Woes

Sunday, March 19, 2017 1:14 PM

Problem(s) Encountered:

From about 4am, CCD readout was abnormal.

(showing pixel value=0 at the 4th quadrant or stripes in the x-axis.)

eg. sample : 170319.0104~0114.fits

Auto dome also stopped from ~ 4 30 am.

I tried  HE box power recycled, restarted prospero.

But these issue still remain.


Replaced the chipset in the dome encoder interface and it appears to be working correctly.  At least for now.

As for the B4K—issue seems to be within the B4K head electronics (HE).  Consulted with Jerry Mason and Rick Pogge.  On the assumption that it was a thermal issue, checked and verified that both fans in the HE were working—they are.  Opened up the side panel and allowed electronics to cool for ~45 minutes, but issue persisted.  Pulled the “junk” board out of the R4K HE and swapped it into the B4K HE.  However issuing a seqinit command led to errors with the sequencer not even being seen.  Verified fiber placement to be correct but problem persisted.  Replaced the B4K junk board and verified that the B4K system still failed in the same manner—it did.  

Consulted the group currently scheduled on the 2.4m and concluded that the best approach at least for tonight was to swap in the R4K and see if they can at least get data.  R4K is on OSMOS and currently cooling.  I will stick around to verify it is working.  Tomorrow, options will be explored as to whether or not to swap to a different spectrograph for the final night of this run.  Of course, Proxy Queue observations start next, so right now, unsure on how we’ll proceed for that.

As an added bit of fun for today, Caliban failed.  Luckily have a spare.  Swapped that in and got it back up and running.  The failed Caliban will be sent back to OSU for repair.